Please read these instructions before posting any event on Fermilab Indico

Indico will be unavailable due to server patching on Wed , June 19th, 7 a.m. to 7:30 a.m. Central Time

ND-GAr: HPgTPC+ECAL Weekly Meeting

US/Central
Disappearance Room (WH13SE) and Zoom

Disappearance Room (WH13SE) and Zoom

    • 1
      Roundtable Updates
      Speakers: Alfons Weber (JG-Uni Mainz & Fermilab), Diego Gonzalez Diaz, Naseem Khan, Tanaz Mohayai (Fermilab)

      Vivek introduced Lauren, a student who is shadowing him.  Welcome!

       

      ECAL

      Vivek:  Leo found that the speed of light incorrect in the MC.  So now seeing less loss than before.  When looking at cluster energy compared to entering track energy, still seeing some large deviations.  Leo says that there is a fhcl file constant that has to be adjusted when the geometry changes (amount of absorber vs amount of scintillator).  This will make E/p = 1.    Presentation maybe in a couple of weeks. 

      Leo:  Working on some MC.  Running into some issues with TTrees and how to copy an entry from one TTree to another and welcomes expert help.

       

      GOAT

      Tanaz: Took a few more chambers out of storage to compare to the one from the Nov/Dec tests that is showing discharge effects.  Visually the new chambers look very clean without any of the discoloration.   With the old chamber do see sparks above ~2200 V in air.    Would like to test one of the chambers from storage in SF6. 

      Alan:  With the chamber showing discharge effects, couldn't see any signs of damage by eye.  But did see sparking in in air in the region near where the source was.  So want to take some of the unused ones and compare to the used one.  Gas breakdown for SF6 in air that is equivalent to 10 atm Ar:CO2

      Tanaz:  Will do these tests in the shipping boxes and will go up to 3000 V.  Eventually also want to take data at 1 atm with a source in argon CO2.

       

      GORG

      Tanaz:  Plans to do first test at 1 atm in a cleanroom.  Need to open box to put source in.  Has identified some options.

      Leo says can store some chambers in Edwards building.

       

      TOAD

      Naseem:  No much hardware updates this week.  Have ordered hydraulic oil and have some power supply options.  Then can pressure test vessel and do full ORC.  Gas lines have been run. Gas on site, but not at FTBF yet.   Having some trouble finding a source.  Kathy says none of the ones on site currently can be used.  Must be able to withstand vacuum and 5 atm, and the sources that FNAL has can't go below 0.25 bar.  Tanaz says need a class 3 source to go to high pressure.    Will maybe look into ordering one.    Have one agregator, one TIP, and one frontendcard at FNAL.  Five of the cards from Pittsburgh are in London and they were checked and all look good.  An order put in for all 156 cards.  Will be tested in UK first and all expected at FNAL by mid-to-late March.  Will try to start with the 5 cards for now.  On software side, has basically finished putting TOAD into DUNE-DAQ, not linked to PACMAN. 

       

      Light collection

      no updates.

       

      Software

      Federico: Validating toy mc and want to check if momentum resolution matches standard formulas for a TPC. Seeing some strange effects, but getting close to giving a longer update.

      Francisco:  Has finished making some changed to ND CAFmaker.  And now can make some files with tracks and ecal info.  Still missing matching to true geant4 particles.